1.a. How long (in ns) does it take light to travel 1.0 m in vacuum?
b. What distance does light travel in water, glass, cubiczirconia during the time it travels in 1.0m vacuum?
Answer #1
distance moved S = 1 m
speed of light in vaccum c = 3 * 10 ^ 8 m / s
So, required time t = S / c
                             = 0.3333 * 10 ^ -8 s
                              =3.333 ns
(b). speed of light in water v = c / n
where n = index of refraction of water = 4 / 3
So, v = 2.25 * 10 ^ 8 m / s
So, distance travel in water S ' = v t
                                              = 0.75 m
(c). speed of light in glass v = c / n
where n = index of refraction of glass = 1.5
So, v = 2 * 10 ^ 8 m / s
So, distance travel in water S ' = v t
                                              = 0.67 m